# Two tales for a geometric Jensen–Shannon divergence

### v1

@@ 2 A novel definition G-JSD extended to positive measures
$D_{\mathrm{JS}}^{A}(p_{1},p_{2})=D_{\mathrm{JS}}^{\tilde{A}}(p_{1},p_{2}).$
However, when the weighted mean $M_{\alpha}$ differs from the weighted arithmetic mean (i.e., $M_{\alpha}\not=A_{\alpha}$), the two definitions of the M-JSD $D_{\mathrm{JS}}^{M}$ and extended M-JSD $D_{\mathrm{JS}}^{\tilde{M}}$ differ by the gap expressed in Eq. 10.
In general, we have $Z_{\alpha}>\log(Z_{\alpha})$ for $Z_{\alpha}>0$, and $Z_{\alpha}-\log(Z_{\alpha})-1<0$ for $Z_{\alpha}<1$ and $Z_{\alpha}-\log(Z_{\alpha})-1\geq 0$ for $Z_{\alpha}\geq 1$.
@@ 2.1 Power JSDs and (extended) min-JSD and max-JSD
Proposition 3
$D_{\mathrm{JS}}^{G}(p_{1},p_{2})=\frac{1}{4}\,D_{J}(p_{1},p_{2})-D_{B}(p_{1},p_{2}),$
where $D_{J}(p_{1},p_{2})=\int(p_{1}-p_{2})\log\frac{p_{1}}{p_{2}}\,{\mathrm{d}\mu}$ is Jeffreys’ divergence [14, 28] and $D_{B}(p_{1},p_{2})=-\log\int\sqrt{p_{1}p_{2}}\,{\mathrm{d}\mu}=-\log Z^{G}(p_{1},p_{2})$ is the Bhattacharrya distance. [^3] The Bhattacharrya distance is not a metric distance as it violates the triangle inequality of metrics.
Proof:
We have:
$D_{\mathrm{JS}}^{G}(p_{1},p_{2})$ $:=$ $\frac{1}{2}\left(D_{\mathrm{KL}}(p_{1},(p_{1}p_{2})^{G})+D_{\mathrm{KL}}(p_{2},(p_{1}p_{2})^{G})\right),$
$=$ $\frac{1}{2}\left(\int\left(p_{1}(x)\log\frac{p_{1}(x)\,Z^{G}(p_{1},p_{2})}{\sqrt{p_{1}(x)p_{2}(x)}}+p_{2}(x)\log\frac{p_{2}(x)\,Z^{G}(p_{1},p_{2})}{\sqrt{p_{1}(x)p_{2}(x)}}\right){\mathrm{d}\mu}(x)\right),$
$=$ $\frac{1}{2}\left(\int\left(p_{1}(x)+p_{2}(x)\right)\log Z^{G}(p_{1},p_{2})\,{\mathrm{d}\mu}(x)+\frac{1}{2}D_{\mathrm{KL}}(p_{1},p_{2})+\frac{1}{2}D_{\mathrm{KL}}(p_{2},p_{1})\right),$
$=$ $\log Z^{G}(p_{1},p_{2})+\frac{1}{4}D_{J}(p_{1},p_{2}),$
$=$ $\frac{1}{4}D_{J}(p_{1},p_{2})-D_{B}(p_{1},p_{2}).$
$\square$
Corollary 1 (G-JSD upper bound)
We have the upper bound $D_{\mathrm{JS}}^{G}(p,q)\leq\frac{1}{4}D_{J}(p,q)$.
Proof:
Since $D_{B}\geq 0$ and $D_{\mathrm{JS}}^{G}(p_{1},p_{2})=\frac{1}{4}D_{J}(p_{1},p_{2})-D_{B}(p_{1},p_{2})$, we have $D_{\mathrm{JS}}^{G}(p,q)\leq\frac{1}{4}D_{J}(p,q)$. $\square$
We have $\lim_{\gamma\rightarrow-\infty}P_{\gamma}(a,b)=\min(a,b)$ and $\lim_{\gamma\rightarrow+\infty}P_{\gamma}(a,b)=\max(a,b)$. Thus we can define both (extented) $\min$-JSD and (extented) $\max$-JSD. Using the fact that $\min(a,b)=\frac{a+b}{2}-\frac{1}{2}|a-b|$ and $\max(a,b)=\frac{a+b}{2}+\frac{1}{2}|a-b|$, we obtain the extremal mixture normalization terms as follows:
$Z_{\min}(p_{1},p_{2})$ $=$ $\int\min(p_{1},p_{2})\,{\mathrm{d}\mu}=1-D_{\mathrm{TV}}(p_{1},p_{2}),$ (12)
$Z_{\max}(p_{1},p_{2})$ $=$ $\int\max(p_{1},p_{2})\,{\mathrm{d}\mu}=1+D_{\mathrm{TV}}(p_{1},p_{2}),$ (13)
where $D_{\mathrm{TV}}(p_{1},p_{2})=\frac{1}{2}\int|p_{1}-p_{2}|\,{\mathrm{d}\mu}$ is the total variation distance.
@@
$0\leq D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})\leq D_{\mathrm{TV}}(p_{1},p_{2}),$ (14)
where $D_{\mathrm{TV}}(p_{1},p_{2})=\frac{1}{2}\,\int|p_{1}-p_{2}|\,{\mathrm{d}\mu}$ denotes the total variation distance. Furthermore, the following identity relates the two types of $\max$-JSDs:
$D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})=D_{\mathrm{JS}}^{\widetilde{\max}}(p_{1},p_{2})+D_{\mathrm{TV}}(p_{1},p_{2})-\log D_{\mathrm{TV}}(p_{1},p_{2}).$ (15)
@@
$D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2}){:=}\frac{1}{2}\,\int\left(p_{1}\log\frac{p_{1}}{\max(p_{1},p_{2})}+p_{2}\log\frac{p_{2}}{\max(p_{1},p_{2})}+2\max(p_{1},p_{2})-(p_{1}+p_{2})\right)\,{\mathrm{d}\mu}.$
Since both $\log\frac{p}{\max(p,q)}\leq 0$ and $\log\frac{q}{\max(p,q)}\leq 0$, and $\max(a,b)=\frac{a+b}{2}+\frac{1}{2}|b-a|$, we have
$D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})\leq\int\left(\frac{p+q}{2}+\frac{1}{2}|q-p|-\frac{p+q}{2}\right)\,{\mathrm{d}\mu}.$
That is, $D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})\leq D_{\mathrm{TV}}(p_{1},p_{2})$ where $D_{\mathrm{TV}}(p_{1},p_{2})$ is the total variation distance.
@@
$=$ $D_{\mathrm{TV}}(p_{1},p_{2})-\log(1+D_{\mathrm{TV}}(p_{1},p_{2}))\geq 0,$
since $0\leq D_{\mathrm{TV}}\leq 1$. Thus $D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})\geq D_{\mathrm{JS}}^{{\max}}(p_{1},p_{2})$. $\square$
@@
$D_{\mathrm{JS}^{+}}^{\widetilde{\min}}(p_{1},p_{2})\geq\frac{1}{4}D_{J}(p_{1},p_{2})-D_{\mathrm{TV}}(p_{1},p_{2}),$
where $D_{J}(p_{1},p_{2}){:=}D_{\mathrm{KL}}(p_{1},p_{2})+D_{\mathrm{KL}}(p_{2},p_{1})=\int(p_{1}-p_{2})\log\frac{p_{1}}{p_{2}}\,{\mathrm{d}\mu}$ is Jeffreys’ divergence [14, 28] and
$D_{\mathrm{JS}^{+}}^{\widetilde{\min}}(p_{1},p_{2})=D_{\mathrm{JS}}^{{\min}}(p_{1},p_{2})-D_{\mathrm{TV}}(p_{1},p_{2})+\log(1-D_{\mathrm{TV}}(p_{1},p_{2})).$
@@
$=$ $-D_{\mathrm{TV}}(p_{1},p_{2})-\log(1-D_{\mathrm{TV}}(p_{1},p_{2}))\geq 0,$
since $0\leq D_{\mathrm{TV}}\leq 1$. Note that the gap can be arbitrarily large when $D_{\mathrm{TV}}(p_{1},p_{2})\rightarrow 1^{-}$.
Thus we have $D_{\mathrm{JS}^{+}}^{\widetilde{\min}}(p_{1},p_{2})\geq D_{\mathrm{JS}}^{{\min}}(p_{1},p_{2})$.
@@
$D_{\mathrm{JS}^{+}}^{\widetilde{\min}}(p_{1},p_{2})$ $=$ $\frac{1}{2}\left(\int(p_{1}\log\frac{p_{1}}{\min(p_{1},p_{2})}+p_{2}\log\frac{p_{2}}{\min(p_{1},p_{2})}+2\min(p_{1},p_{2})-(p_{1}+p_{2})\right)\,{\mathrm{d}\mu},$
@@
  • When $\sigma_{1}=\sigma_{2}=\sigma_{2}$, we have
    $\mathrm{TV}(p_{1},p_{2})=\frac{1}{2}\left|\Phi(x^{*};\mu_{2},\sigma_{2})-\Phi(x^{*};\mu_{1},\sigma_{1})\right|,$ (16)
@@ 3.1 Exponential families
$F_{\theta}(\theta)=\frac{1}{2}\left(d\log\pi-\log|\theta_{M}|+\frac{1}{2}\theta_{v}^{\top}\theta_{M}^{-1}\theta_{v}\right),$ (27)
@@
$\theta=(\theta_{v},\theta_{M})=\left\{\begin{array}[]{ll}\theta_{v}(\lambda)=\lambda_{M}^{-1}\lambda_{v}=\Sigma^{-1}\mu\\ \theta_{M}(\lambda)=\frac{1}{2}\lambda_{M}^{-1}=\frac{1}{2}\Sigma^{-1}\end{array}\right$. $\Leftrightarrow$ $\lambda=(\lambda_{v},\lambda_{M})=\left\{\begin{array}[]{ll}\lambda_{v}(\theta)=\frac{1}{2}\theta_{M}^{-1}\theta_{v}=\mu\\ \lambda_{M}(\theta)=\frac{1}{2}\theta_{M}^{-1}=\Sigma\end{array}\right$. (34)
@@
$D_{\mathrm{JS}^{+}}^{\tilde{G}_{\alpha},\beta}(p_{\mu_{1},\Sigma_{1}},p_{\mu_{2},\Sigma_{2}})\leq D_{\mathrm{JS}}^{{G}_{\alpha},\beta}(p_{\mu_{1},\Sigma_{1}},p_{\mu_{2},\Sigma_{2}}).$
@@
$D_{\mathrm{JS}}^{G}(p_{\theta_{1}},p_{\theta_{2}})={\frac{1}{4}(\theta_{2}-\theta_{1})^{\top}(\nabla F(\theta_{2})-\nabla F(\theta_{1}))}-{\left(\frac{F(\theta_{1})+F(\theta_{2})}{2}-F\left(\frac{\theta_{1}+\theta_{2}}{2}\right)\right)},$ (39)
since $D_{J}(p_{\theta_{1}},p_{\theta_{2}})=\left\langle\theta_{2}-\theta_{1},\nabla F(\theta_{2})-\nabla F(\theta_{1})\right\rangle$ amounts to a symmetrized Bregman divergence [28].
@@
We have $D_{J}(p_{\theta_{1}},p_{\theta_{2}})=(\theta_{2}-\theta_{1})^{\top}(\nabla F(\theta_{2})-\nabla F(\theta_{1}))$ and $D_{J}(p_{\theta_{1}},p_{\theta_{2}})=J_{F}(\theta_{1},\theta_{2})$. $\square$
@@
$D_{\mathrm{JS}}^{G}(p_{\theta_{1}},p_{\theta_{2}})$ $=$ ${\frac{1}{4}(\theta_{2}-\theta_{1})^{\top}(\nabla F(\theta_{2})-\nabla F(\theta_{1}))}-{\left(\frac{F(\theta_{1})+F(\theta_{2})}{2}-F\left(\frac{\theta_{1}+\theta_{2}}{2}\right)\right)},$ (40)
$D_{\mathrm{JS}}^{\tilde{G}}(p_{\theta_{1}},p_{\theta_{2}})$ $=$ $\frac{1}{4}\left\langle\theta_{2}-\theta_{1},\nabla F(\theta_{2})-\nabla F(\theta_{1})\right\rangle-\exp(-J_{F}(\theta_{1},\theta_{2}))-1,$ (41)
$D_{\mathrm{JS}^{*}}^{G}(p_{\theta_{1}},p_{\theta_{2}})$ $=$ $J_{F}(\theta_{1},\theta_{2})$ (42)

### v3

@@ 2.1 Definition and properties
$\mathrm{JS}_{A}(p_{1},p_{2})={\mathrm{JS}}_{\tilde{A}}(p_{1},p_{2}).$
However, when the weighted mean $M_{\alpha}$ differs from the weighted arithmetic mean (i.e., $M_{\alpha}\not=A_{\alpha}$), the two definitions of the M-JSD $\mathrm{JS}_{M}$ and extended M-JSD ${\mathrm{JS}}_{\tilde{M}}$ differ by the gap expressed in Eq. 17.
Remark 8
When information is measured in bits, logarithms are taken to base $2$ and when information is measured in nats, base $e$ is considered. Thus we shall generally consider the gap $\Delta_{b}=Z_{\alpha}-\log_{b}(Z_{\alpha})-1$ where $b$ denotes the base of the logarithm. When $b=e$, we have $\Delta_{e}\geq 0$ for all $Z_{\alpha}>0$. When $b=2$, we have $\Delta_{2}=Z_{\alpha}-\log_{2}(Z_{\alpha})-1\geq 0$ when $0<Z_{\alpha}\leq 1$ or $Z_{\alpha}\geq 2$. But since $Z_{\alpha}\leq 2$ (see Eq. 7), the condition simplifies to $\Delta_{2}\geq 0$ if and only if $Z_{\alpha}\leq 1$.
Remark 9
Although $\sqrt{\mathrm{JS}}$ is a metric distance [18], $\sqrt{\mathrm{JS}_{G}}$ is not a metric distance as the triangle inequality is not satisfied. It suffices to report a counterexample of the triangle inequality for a triple of points $p_{1}$, $p_{2}$, and $p_{3}$: Consider $p_{1}=(0.55,0.45)$, $p_{2}=(0.002,0.998)$, and $p_{3}=(0.045,0.955)$. Then we have $\sqrt{\mathrm{JS}_{G}}(p_{1},p_{2})\approx 1.0263227\ldots$, $\sqrt{\mathrm{JS}_{G}}(p_{1},p_{3})\approx 0.63852342\ldots$, and $\sqrt{\mathrm{JS}_{G}}(p_{3},p_{2})\approx 0.19794622\ldots$. The triangle inequality fails with an error of
$\sqrt{\mathrm{JS}_{G}}(p_{1},p_{2})-(\sqrt{\mathrm{JS}_{G}}(p_{1},p_{3})+\sqrt{\mathrm{JS}_{G}}(p_{3},p_{2}))\approx 0.1898531\dots.$
Similarly, the triangle inequality also fails for the extended G-JSD: We have $\sqrt{\mathrm{JS}^{+}_{G}(p_{1},p_{2})}\approx 1.0788275\dots$, $\sqrt{\mathrm{JS}^{+}_{G}(p_{1},p_{3})}\approx 0.6691922\ldots$, and $\sqrt{\mathrm{JS}^{+}_{G}(p_{3},p_{2})}\approx 0.1984633\ldots$ with a triangle inequality defect value of
$\sqrt{\mathrm{JS}^{+}_{G}(p_{1},p_{2})}-(\sqrt{\mathrm{JS}^{+}_{G}(p_{1},p_{3})}+\sqrt{\mathrm{JS}^{+}_{G}(p_{3},p_{2})})\approx 0.2111719\ldots.$
@@ 2.2 Power JSDs and (extended) min-JSD and max-JSD
Let $P_{\gamma}(a,b)=P_{\gamma,\frac{1}{2}}(a,b)$. We have $\lim_{\gamma\rightarrow-\infty}P_{\gamma}(a,b)=\min(a,b)$ and $\lim_{\gamma\rightarrow+\infty}P_{\gamma}(a,b)=\max(a,b)$. Thus we can define both (extented) $\min$-JSD and (extented) $\max$-JSD. Using the fact that $\min(a,b)=\frac{a+b}{2}-\frac{1}{2}|a-b|$ and $\max(a,b)=\frac{a+b}{2}+\frac{1}{2}|a-b|$, we obtain the extremal mixture normalization terms as follows:
$Z_{\min}(p_{1},p_{2})$ $=$ $\int\min(p_{1},p_{2})\,{\mathrm{d}\mu}=1-\mathrm{TV}(p_{1},p_{2}),$ (19)
$Z_{\max}(p_{1},p_{2})$ $=$ $\int\max(p_{1},p_{2})\,{\mathrm{d}\mu}=1+\mathrm{TV}(p_{1},p_{2}),$ (20)
where $\mathrm{TV}(p_{1},p_{2})=\frac{1}{2}\int|p_{1}-p_{2}|\,{\mathrm{d}\mu}$ is the total variation distance.
@@
$0\leq{\mathrm{JS}^{+}}_{\widetilde{\max}}(p_{1},p_{2})\leq\mathrm{TV}(p_{1},p_{2}).$ (21)
Furthermore, the following identity relates the two types of $\max$-JSDs:
${\mathrm{JS}^{+}}_{\widetilde{\max}}(p_{1},p_{2})={\mathrm{JS}}_{\widetilde{\max}}(p_{1},p_{2})+\mathrm{TV}(p_{1},p_{2})-\log\left(1+\mathrm{TV}(p_{1},p_{2})\right).$ (22)
@@
${\mathrm{JS}^{+}}_{\widetilde{\max}}(p_{1},p_{2}){:=}\frac{1}{2}\,\int\left(p_{1}\log\frac{p_{1}}{\max(p_{1},p_{2})}+p_{2}\log\frac{p_{2}}{\max(p_{1},p_{2})}+2\max(p_{1},p_{2})-(p_{1}+p_{2})\right)\,{\mathrm{d}\mu}.$
Since both $\log\frac{p_{1}}{\max(p_{1},p_{2})}\leq 0$ and $\log\frac{p_{2}}{\max(p_{1},p_{2})}\leq 0$, and $\max(a,b)=\frac{a+b}{2}+\frac{1}{2}|b-a|$, we have
${\mathrm{JS}^{+}}_{\widetilde{\max}}(p_{1},p_{2})\leq\int\left(\frac{p_{1}+p_{2}}{2}+\frac{1}{2}|p_{2}-p_{1}|-\frac{p_{1}+p_{2}}{2}\right)\,{\mathrm{d}\mu}.$
That is, ${\mathrm{JS}^{+}}_{\widetilde{\max}}(p_{1},p_{2})\leq\mathrm{TV}(p_{1},p_{2})$.
@@
$=$ $\mathrm{TV}(p_{1},p_{2})-\log(1+\mathrm{TV}(p_{1},p_{2}))\geq 0,$
since $0\leq\mathrm{TV}\leq 1$. Thus ${\mathrm{JS}^{+}}_{\widetilde{\max}}(p_{1},p_{2})\geq{\mathrm{JS}}_{{\max}}(p_{1},p_{2})$. $\square$
@@
${\mathrm{JS}^{+}}_{\widetilde{\min}}(p_{1},p_{2})\geq\frac{1}{4}\,J(p_{1},p_{2})-\mathrm{TV}(p_{1},p_{2}),$
where $J(p_{1},p_{2}){:=}{\mathrm{KL}}(p_{1},p_{2})+{\mathrm{KL}}(p_{2},p_{1})=\int(p_{1}-p_{2})\log\frac{p_{1}}{p_{2}}\,{\mathrm{d}\mu}$ is Jeffreys’ divergence [25] and
${\mathrm{JS}^{+}}_{\widetilde{\min}}(p_{1},p_{2})={\mathrm{JS}}_{{\min}}(p_{1},p_{2})-\mathrm{TV}(p_{1},p_{2})+\log(1-\mathrm{TV}(p_{1},p_{2})).$
@@
$=$ $-\mathrm{TV}(p_{1},p_{2})-\log(1-\mathrm{TV}(p_{1},p_{2}))\geq 0,$
since $x-\log(1-x)\geq 0$ for $x\leq 1$. Note that the gap can be arbitrarily large when $\mathrm{TV}(p_{1},p_{2})\rightarrow 1^{-}$.
Thus we have ${\mathrm{JS}^{+}}_{\widetilde{\min}}(p_{1},p_{2})\geq{\mathrm{JS}}_{{\min}}(p_{1},p_{2})$.
@@
${\mathrm{JS}^{+}}_{\widetilde{\min}}(p_{1},p_{2})$ $=$ $\frac{1}{2}\left(\int(p_{1}\log\frac{p_{1}}{\min(p_{1},p_{2})}+p_{2}\log\frac{p_{2}}{\min(p_{1},p_{2})}+2\min(p_{1},p_{2})-(p_{1}+p_{2})\right)\,{\mathrm{d}\mu},$
@@
  • When $\sigma_{1}=\sigma_{2}=\sigma$, we have
    $\mathrm{TV}(p_{1},p_{2})=\frac{1}{2}\left|\Phi(x^{*};\mu_{2},\sigma)-\Phi(x^{*};\mu_{1},\sigma)\right|,$ (23)
@@ 3.1 Exponential families
$F_{\theta}(\theta){:=}\frac{1}{2}\left(d\log\pi-\log|\theta_{M}|+\frac{1}{2}\theta_{v}^{\top}\theta_{M}^{-1}\theta_{v}\right),$ (34)
@@
$\theta=(\theta_{v},\theta_{M})=\left\{\begin{array}[]{ll}\theta_{v}(\lambda)=\lambda_{M}^{-1}\lambda_{v}=\Sigma^{-1}\mu\\ \theta_{M}(\lambda)=\frac{1}{2}\lambda_{M}^{-1}=\frac{1}{2}\Sigma^{-1}\end{array}\right$. $\Leftrightarrow$ $\lambda=(\lambda_{v},\lambda_{M})=\left\{\begin{array}[]{ll}\lambda_{v}(\theta)=\frac{1}{2}\theta_{M}^{-1}\theta_{v}=\mu\\ \lambda_{M}(\theta)=\frac{1}{2}\theta_{M}^{-1}=\Sigma\end{array}\right$.
@@
${\mathrm{JS}^{+}}_{\tilde{G}_{\alpha},\beta}(p_{\mu_{1},\Sigma_{1}},p_{\mu_{2},\Sigma_{2}})\leq\mathrm{JS}_{{G}_{\alpha},\beta}(p_{\mu_{1},\Sigma_{1}},p_{\mu_{2},\Sigma_{2}}).$
@@
$\mathrm{JS}_{G}(p_{\theta_{1}},p_{\theta_{2}})={\frac{1}{4}(\theta_{2}-\theta_{1})^{\top}(\nabla F(\theta_{2})-\nabla F(\theta_{1}))}-{\left(\frac{F(\theta_{1})+F(\theta_{2})}{2}-F\left(\frac{\theta_{1}+\theta_{2}}{2}\right)\right)},$ (41)
since $J(p_{\theta_{1}},p_{\theta_{2}})=\left\langle\theta_{2}-\theta_{1},\nabla F(\theta_{2})-\nabla F(\theta_{1})\right\rangle$ amounts to a symmetrized Bregman divergence.
@@
We have $J(p_{\theta_{1}},p_{\theta_{2}})=(\theta_{2}-\theta_{1})^{\top}(\nabla F(\theta_{2})-\nabla F(\theta_{1}))$ and $J(p_{\theta_{1}},p_{\theta_{2}})=J_{F}(\theta_{1},\theta_{2})$. $\square$
@@
$\mathrm{JS}_{G}(p_{\theta_{1}},p_{\theta_{2}})$ $=$ ${\frac{1}{4}(\theta_{2}-\theta_{1})^{\top}(\nabla F(\theta_{2})-\nabla F(\theta_{1}))}-{\left(\frac{F(\theta_{1})+F(\theta_{2})}{2}-F\left(\frac{\theta_{1}+\theta_{2}}{2}\right)\right)},$ (42)
$\mathrm{JS}_{\tilde{G}}(p_{\theta_{1}},p_{\theta_{2}})$ $=$ $\frac{1}{4}\left\langle\theta_{2}-\theta_{1},\nabla F(\theta_{2})-\nabla F(\theta_{1})\right\rangle-\exp(-J_{F}(\theta_{1},\theta_{2}))-1,$ (43)
${\mathrm{JS}^{*}}_{G}(p_{\theta_{1}},p_{\theta_{2}})$ $=$ $J_{F}(\theta_{1},\theta_{2})$ (44)
Remark 12
Given two densities $p_{1}$ and $p_{2}$, the family $\mathcal{G}$ of geometric mixtures $\{(p_{1}p_{2})_{G_{\alpha}}\propto p_{1}^{\alpha}\,p_{2}^{1-\alpha}{\ :\ }\alpha\in(0,1)\}$ forms a 1D exponential family that has been termed likelihood ratio exponential family [22] (LREF). The cumulant function of this LREF is $F(\alpha)=-B_{\alpha}(p_{1},p_{2})$. Hence, $\mathcal{G}$ has also been called a Bhattacharyya arc or Hellinger arc in the literature [10]. However, notice that ${\mathrm{KL}}(p_{i}:(p_{1}p_{2})_{G_{\alpha}})$ does not amount necessarily to a Bregman divergence because neither $p_{1}$ nor $p_{2}$ belongs to $\mathcal{G}$.